### v1

# 22-large sets are sets of Bohr recurrence

Abstract
Let $\alpha_{1},\cdots,\alpha_{d}$ be real numbers, and let $S$ be the set of integers $s$ so that $\alpha_{i}s||_{\mathbb{R}/\mathbb{Z}}>\delta$ for some $i$ and some fixed $\delta>0$. We prove $S$ is not 22-large”, i.e. there is a $2$-coloring of $\mathbb{N}$ that avoids arbitrarily long arithmetic progressions with common differences in $S$.
1 Introduction
One of the oldest and most classic results in Ramsey theorem is van der Waerden’s theorem, that any finite coloring of $\mathbb{N}$ has arbitrarily long arithmetic progressions in some color [11].
A natural question, with deep connections to dynamics, is what happens if the common differences are restricted to be in some set $S$. In the 1990s, Brown, Graham, and Landman [3] defined a set $S$ to be “large” if van der Waerden holds with restricted differences in the set $S$.
Definition 1.1 (Large Sets [3]).
We say that $S\subset\mathbb{N}$ is “large” if whenever $\mathbb{N}$ is finitely colored, there are arbitrarily long arithmetic progressions with common differences in $S$.
A large set is also referred to as a set of multiple recurrence or a set of topological multiple recurrence in the dynamics literature. See [1] for discussion about the equivalence.
Brown, Graham, and Landman also defined “2-large”, an analogous notion for $2$-colorings.
Definition 1.2 (2-Large Sets [3]).
We say that $S\subset\mathbb{N}$ is “large” if whenever $\mathbb{N}$ is two-colored, there are arbitrarily long arithmetic progressions with common differences in $S$.
It is trivial to see that large sets are $2$-large sets. Brown, Graham, and Landman conjectured that the reverse is also true. This is referred to as the “Large Sets Conjecture” in the literature.
Conjecture 1.3 (Large Sets Conjecture [3]).
If $S$ is $2$-large, then $S$ is large.
While their paper was published in 1999, the Open Problem Garden claims [4] that the conjecture was formulated in 1995. Shortly after the original formulation of the large sets conjecture, Bergelson and Leibman [2] proved that if $S$ is the values of an integer polynomial, then $S$ is large. In particular, this is a special case of the polynomial van der Waerden theorem of Bergelson and Leibman.
Perhaps it is easier to think about which sets are not large. We must determine if all such sets are $2$-large. But what do non-large sets look like? This is in fact the main subject of the paper of Brown, Graham, and Landman [3]. Until recently the only known examples of non-large sets were sets that were not sets of “nil-Bohr recurrence”, which are known to be non-large, and [10] asked whether or not these are the only such sets. Very recently, the author gave a new example of a non-large set [1].
Let us start with a basic example of a non-large set. If $S$ consists only of numbers that are far away from a multiple of some $\alpha$, then $S$ is not large, because we can color $x$ by $x$ modulo $\alpha$ up to a small error $\varepsilon$ and if $x,x+d$ are the same color then their difference $d$ will be very close to a multiple of $\alpha$. It is known (see [6] and Theorem 7.4 of [9]) that such sets are not $2$-large).
Definition 1.4 (One-dimensional Bohr recurrence).
A set $S$ is a set of one-dimensional Bohr recurrence if for any $\alpha$, there exists $s\in S$ so that $s\alpha||_{\mathbb{R}/\mathbb{Z}}<\varepsilon$.
In other words, it is known that if $S$ is $2$-large, then it is a set of one-dimensional Bohr recurrence.
Definition 1.5 (Bohr recurrence).
A set $S$ is a set of ($d$-dimensional) Bohr recurrence if for any $\alpha_{1},\cdots,\alpha_{d},\varepsilon$, there exists $s\in S$ so that $s\alpha_{i}||_{\mathbb{R}/\mathbb{Z}}<\varepsilon$ for $1\leq i\leq d$.
If a set $S$ is not a set of Bohr recurrence, then it is readily seen not to be large. One can color $n$ by taking the map $n\to(\{\alpha_{1}n\},\cdots,\{\alpha_{d}n\})\in(\mathbb{R}/\mathbb{Z})^{d}$, partitioning $(\mathbb{R}/\mathbb{Z})^{d}$ into $\varepsilon$-balls, and coloring $n$ by which $\varepsilon$-ball its image lands in. More generally, if $S$ is not a set of “nil-Bohr recurrence”, then it is not large. We are ready to state our main result.
Theorem 1.6 (Main Theorem).
If there exists $\alpha_{1},\cdots,\alpha_{d}$ so that for all $s\in S$ there exists $1\leq i\leq d$ so that $s\alpha_{i}||_{\mathbb{R}/\mathbb{Z}}>\varepsilon$ (i.e. $S$ is not a set of Bohr recurrence), then $S$ is not $2$-large.
We also state the contrapositive, because it is a bit cleaner.
Theorem 1.7 (Main Theorem, contrapositive).
If $S$ is $2$-large, then it is a set of Bohr recurrence.
This result generalizes many of the examples from [3]. In particular, [3] presents many examples of non-large sets which are in fact not Bohr recurrent, such as lacunary sets.
While we will not formally define “nil-Bohr recurrence” in the paper, essentially it is the higher-degree generalization of Bohr recurrence, so for instance if $s^{2}$ is far from a multiple of $\alpha$ for $s\in S$ then the identity $(x+2s)^{2}-2(x+s)^{2}+x^{2}=2s^{2}$ gives an obstruction to largeness. One can replace $s^{2}$ with any higher degree polynomial or “generalized polynomial” involving floor functions, see [1] for details.
In the next two sections, we only deal with $2$-largeness and not largeness, so we use the word “large” in the usual way.
2 Coloring and Proof Outline
Before describing the proof of Theorem 1.6, we give a remark about the technique from [6] and [9] and why it will not generalize, and we explain what we need to do instead. The proof of [6] and [9] proceeds by considering a map $f$ from $\mathbb{Z}$ to $\mathbb{R}/\mathbb{Z}$ taking $n\to\{n\alpha\}$, and coloring $\mathbb{R}/\mathbb{Z}$ in two colors. In particular, if one makes the interval $[0,1/2)$ is red and the interval $[1/2,1)$ is blue then the length of a monochromatic arithmetic progression whose common difference is bounded away from 0 in $\mathbb{R}/\mathbb{Z}$ will be bounded. This does not generalize.
Theorem 2.1.
Let $d$ be a large positive integer. Then a two-coloring of $(\mathbb{R}/\mathbb{Z})^{d}$ will contain an infinitely long arithmetic progressions with nonzero common difference $s$, so that all coordinates of $s\in(\mathbb{R}/\mathbb{Z})^{d}$ are 0 or $1/3$.
Proof.
Look only at the points $\{0,1/3,2/3\}^{d}$. By Hales-Jewett we can find a combinatorial line, which correspondents to a three-term and thus an infinitely long arithmetic progression. ∎
Instead we need to take the universal cover of the torus $(\mathbb{R}/\mathbb{Z})^{d}$ and consider the map from $f:\mathbb{Z}\to\mathbb{R}^{d}$ given by $x\to(\alpha_{1}x,\cdots,\alpha_{d}x)$. Then we color $\mathbb{R}^{d}$ in the following way. Pick integers $N_{1}\gg N_{0}$ (by this we mean that $N_{1}$ is a sufficiently large function of $N_{0}$) chosen with foresight so that $k:=\frac{N_{1}}{N_{0}}$ and $N_{0}$ are divisible by a sufficiently large factorial (depending on $\delta$). We will have the coloring be periodic mod $N_{1}$ in all directions. So then it suffices to consider a coloring of $(\mathbb{R}/N_{1}\mathbb{Z})^{d}$, or $0,N_{1})^{d}$. This object is naturally decomposed into $k^{d}$ translates of $0,N_{0})^{d}$, called cells. We will color each cell red or blue randomly. We refer to this object, the $k^{d}$ cells all colored randomly, as the “chessboard”. We can identify all of $\mathbb{R}^{d}$ and thus all of $\mathbb{Z}$ with this chessboard, through pulling back the maps $\mathbb{Z}\to\mathbb{R}^{d}\to(\mathbb{R}/N_{1}\mathbb{Z})^{d}$.
Then, if we take $N_{2}\gg N_{1}$ (of course it will be much bigger than any function of $\delta$), any arithmetic progression with a common difference $x$ of length $N_{2}$ so that $\alpha_{i}x||>\delta$ will meet both red and blue squares. The idea is that it will trace out a line in the “chessboard”. There are only $k^{O_{d}(1)}$ many such “lines” and each has a probability of $2^{-\Omega(k)}$ of being monochromatic. So if $k$ is chosen large enough, a union bound gives us what we want by a simple probabilistic argument. In some sense, this is a combinatorial phrasing of the fact that a torus rotation has zero entropy.
To show this, we must use that the arithmetic progression will return near its original position in the chessboard within $s=O(k^{d})$ steps. If it does not go too close to its original position in $s$ steps, then going another $s$ steps, and another $s$ steps and so on will cause it to trace out a “line” in far fewer than $N_{2}$ steps. If it is very close to the original position in $s$ steps, then we finish by a divisibility argument. In particular, we will use the fact that $s\mid N_{1}$. The step size in the chessboard must be very close to a multiple of $N_{1}/s$ chessboard cells in each direction which contradicts that $\alpha_{i}x||>\delta$ for some $i$. The full proof is not much longer.
3 The Proof
Let $\alpha_{1},\cdots,\alpha_{d}\in\mathbb{R}$, and let $S$ be the set of $x$ so that $\alpha_{i}x||>\delta$ for some $i$ and some fixed $\delta>0$. We will prove $S$ is not $2$-large by exhibiting a $2$-coloring.
Define some sufficiently large integers $N_{0}\gg k\gg C\gg\max(d,1/\delta)$, so that $N_{1}=kN_{0}$ is divisible by all integers up to $CN_{1}/N_{0}=Ck$. We can ensure this by for instance setting $N_{0}$ to be a sufficiently large factorial function. We will pick $N_{2}$ which is a sufficiently large as a function of $N_{1}$ and then we will show that there is a coloring so that every $N_{2}$-AP with common differences in $S$ receives both colors.
Consider the map $f:\mathbb{Z}\to\mathbb{R}^{d}$ sending $x\to(\alpha_{1}x,\cdots,\alpha_{d}x)$ and the usual quotient map $g:\mathbb{R}^{d}\to(\mathbb{R}/N_{1}\mathbb{Z})^{d}$. Composing these two maps, we get a map $F:\mathbb{Z}\to(\mathbb{R}/N_{1}\mathbb{Z})^{d}$. We will color the torus $(\mathbb{R}/N_{1}\mathbb{Z})^{d}$ red and blue, and by taking the pre-image under $F$ this gives a coloring of $\mathbb{Z}$. In order to color this torus, which has fundamental domain $0,N_{1})^{d}$ in $(\mathbb{R}/N_{1}\mathbb{Z})^{d}$, we decompose into $(N_{1}/N_{0})^{d}$ copies of $0,N_{0})^{d}$ in the usual way. Each copy will be colored red or blue, uniformly and independently. We claim that with high probability this coloring works, and any arithmetic progression of length $N_{2}$ with common difference in $S$ will have both red and blue elements. We call the object $(\mathbb{R}/N_{1}\mathbb{Z})^{d}$ the $N_{1}$-torus” and we say it is decomposed into $(N_{1}/N_{0})^{d}$$N_{0}$-cells”. Furthermore, we split each $N_{0}$-cell into $C^{d}$ “mini-cells” in the obvious way.
Say we have an arithmetic progression in $\mathbb{Z}$ of the form $x,x+y,\cdots,x+N_{2}y$ with $y\in S$. We take its image under $F$ in the torus $(\mathbb{R}/N_{1}\mathbb{Z})^{d}$. By pigeonhole, at least two of $F(x),F(x+y),\cdots,F(x+(CN_{1}/N_{0})^{d})$ lie in the same mini-cell, say $F(x+ay)$ and $F(x+by)$. This means for some $t=|a-b|\leq(CN_{1}/N_{0})^{d}\ll N_{2}$, we have that $F(x+ty)$ and $F(x)$ are close in the $N_{1}$-torus, differing by at most $N_{0}/C$ in each coordinate.
Since we chose $t\mid N_{2}$, we may consider the sub-progression $x,x+ty,\cdots,x+N_{2}y$ of length $N_{2}/t+1\gg N_{1}$. We first show that $F(ty)$ has some coordinate bounded away from 0 in the $N_{1}$-torus, via a divisibility argument. Let $y^{\prime}\in[0,N_{1})^{d}$ be such that $F(y)\equiv y^{\prime}$ in $(\mathbb{R}/N_{1}\mathbb{Z})^{d}$. Then we have $ty^{\prime}\in[0,tN_{1})^{d}$, and by assumption $y^{\prime}$ is at least $\delta$ away from an integer point in the $L^{\infty}$ distance. Say $ty^{\prime}$ has distance $\eta$ from some point $(b_{1}N_{1},\cdots,b_{d}N_{1})$ where $b_{i}$ are integers, so that $F(ty)$ is $\eta$-close to 0 in $(\mathbb{R}/N_{1}\mathbb{Z})^{d}$ in $L^{\infty}$ distance. Then $y^{\prime}$ has distance at most $\eta/t\leq\eta$ from the integer point $(b_{1}\frac{N_{1}}{t},\cdots,b_{d}\frac{N_{1}}{t})$, and so $\eta\geq\delta$.
Now, we have that $F(ty)$ has some coordinate at least $\delta$-far from 0 in the $N_{1}$-torus, say the $i$th coordinate, and all its coordinates are at most $(N_{1}/CN_{0})^{d}$ in magnitude. Since $N_{2}\gg N_{1}$ is sufficiently large, this means that $F(x),F(x+ty),\cdots F(x+(N_{2}/N_{1})y)$ will all lie on a line in the $N_{1}$-torus.
Every line will meet $\Omega_{d,1/\delta}(k)$ such cells. One can see this by for instance projecting onto the $i$th coordinate. In fact, we can find $k^{O_{d,1/\delta}(1)}$ choices of $\Omega_{d,1/\delta}(k)$ cells (which we call an orbit) so that every line meets one of them, for instance by controlling the angle of the line up to a $1/k^{O_{d,1/\delta}(1)}$ error and the intercept up to a $N_{1}/k^{O_{d,1/\delta}(1)}$ error.
Figure 1: An example of an orbit. On the $N_{1}$-torus, this line wraps around the board.
Hence, there are only $(N_{1}/N_{0})^{O_{d,1/delta}(1)}=k^{O_{d,1/\delta}(1)}$ possible orbits, and in a random coloring each one has a chance exponentially small in $k$ of being monochromatic. So if $k$ is chosen large enough, then we will not have length $N_{2}$ arithmetic progressions that are monochromatic.
4 Further Directions
In general, a set is not “large” in the sense of [3] if it is not a set of multiple recurrence for some dynamical system. Toral rotations are the first example of a dynamical system. One could ask also about increasingly more general classes of dynamical systems such as nilsystems, skew products, and distal systems. Since every dynamical system can be built from distal systems and weak mixing systems by the Furstenberg-Zimmer structure theorem ([7] [13]), this is perhaps a potential approach to settling the $2$-large versus large conjecture in the positive.
We know how to prove this for some nilrotations. If $\alpha$ is irrational, then the set of $n$ for which the fractional part $\{n^{2}\alpha\}$ is bounded away from 0 and $1$, so the set of $n$ so that $\{n^{2}\alpha\}||_{\mathbb{R}/\mathbb{Z}}>\delta$ for some $\delta>0$ can be shown to not be $2$-large by a similar argument, using standard facts about the equidistribution of $n^{2}\alpha$ (see [5], [12]). We strongly suspect that the same holds for all nilsystems. The construction of [1] can be rephrased as coming from a skew product on $(\mathbb{R}/\mathbb{Z})^{5}$. We generally suspect that examples like the one in [1] are not $2$-large as well. However, we do not have strong opinions on what happens for all dynamical systems, i.e. on whether the $2$-large versus large conjecture of [3] is true.
5 Acknowledgments
The author thanks Sohail Farhangi for helpful comments, and thanks GPT-5 for LaTeX help.